If the desired fiscal stimulus is $40 billion and the desired AD increase is $50 billion, we can conclude about the MPC and multiplier?
1
Expert's answer
2018-04-04T07:56:08-0400
MPC = fiscal stimulus / AD = 40/50 = 0.8
multiplier = AD / fiscal stimulus = 50/40 = 1.25